6.B. Find: charge on each capacitor? voltage across each capacitor? C = 3.0uF; C. = 8.0uF: C = 4.0uF: V = 5.0V
Answer #1

C₂ and Cz o re in their equivalent parallel combination Capacitance =C2+ C3 = 8+4 =12 MF 1 C is in series with 12 uf C = 7 3x